OpenStudy (anonymous):

A hybrid car can travel 40 miles per gallon. Approximately how many gallons of fuel will the car need to travel 20 km? [1 mile = 1.6 km] 0.3 0.8 2.5 3.2

OpenStudy (anonymous):

20 Km is 12.4~ miles. If you get 40 miles per gallon. And you need to travel 12.4 miles you can already assume you're traveling less than 1 mile so you can eliminate 2.5 and 3.2. Divide 12.4 by 40 and that should give you the desired answer.

OpenStudy (shreenithaa):

20/1.6=

OpenStudy (shreenithaa):

thats the answer

OpenStudy (anonymous):

That gives you the miles traveled, Now it wants you to take the known that you travel 40 miles on 1 gallon of gas. so 20/1.6 gives you 12.5. where 12.5 is equal to the miles traveled. If you 1 gallon gets you 40 miles then how many gallons for 12.5 miles. Set that as a fraction of 1/40 * x/12.5 Cross multiply and you get 12.5=40x divide both sides you get x=12.5/40 where x=gallons of gas.